8-th International Workshop on Complex structures and vector fields

Sofia, 21-26 August 2006

List of participants

August 18, 2006

LIST of PARTICIPANTS

Toshiaki ADACHI Department of Mathematics, Nagoya Institute of Technology Nagoya 466-8555, Japan e-mail: adachi@nitech.ac.jp

Vestislav Apostolov UQAM Université du Québec Montréal, Case postale 8888, succursale Centre-ville Montréal (Québec), H3C 3P8 Canada e-mail: apostolo@math.uqam.ca

Lilia N. Apostolova Institute of mathematics and informatics, acad. G. Bontchev str., 1113 Sofia, Bulgaria e-mail: liliana@math.bas.bg

Vladimir Balan University Politehnica of Bucharest Faculty of Applied Sciences, Department of Mathematics I, RO-060042, Bucharest, Romania. e-mail: vbalan@mathem.pub.ro

Rossen Dandoloff LPTM, Université de Cergy-Pontoise, F-95302 Cergy-Pontoise Cedex, France e-mail: rossen.dandoloff@u-cergy.fr

Stancho Dimiev Institute of mathematics and informatics, acad. G. Bontchev str., 1113 Sofia, Bulgaria e-mail: sdimiev@math.bas.bg

Stoil Donev Institute for Nuclear Research and Nuclear Energy, 72 blvd.Tzarigradsko chaussee 1784 Sofia, Bulgaria e-mails: sdonev@inrne.bas.bg

Norio Ejiri Department of Mathematics, Meijo University, Tempaku, Nagoya, 468-8502 Japan e-mail: ejiri@ccmfs.meijo-u.ac.jp

Petia Furlinska Plovdiv, Bulgaria e-mail: petiaf@abv.bg

Vladimir Georgiev Dipartimento di Matematica, Università di Pisa, Largo Bruno Pontecorvo 5, 56127 Pisa, Italy e-mail: georgiev@dm.unipi.it

Vladimir Gerdjikov Institute for Nuclear Research and Nuclear Energy, 72 Tzarigradsko chaussee , 1784 Sofia, Bulgaria e-mail: gerjikov@inrne.bas.bg

Georgi Grahovski Institute for Nuclear Research and Nuclear Energy, 72 Tzarigradsko chaussee, 1784 Sofia, Bulgaria e-mail: grah@inrne.bas.bg

Kostadin Gribachev Faculty of Mathematics and Informatics, University of Plovdiv, Plovdiv 4003, Bulgaria

Dobrinka K. Gribacheva-Shtarbeva Faculty of Mathematics and Informatics, University of Plovdiv, Plovdiv 4003, Bulgaria e-mail: dobrinka@pu.acad.bg

Hideya Hashimoto Meijo University, Tempaku, Nagoya, Japan e-mail: hhashi@ccmfs.meijo-u.ac.jp

Milen Hristov University of Veliko Tarnovo, 2 T. Tarnovski str., 5000 Veliko Tarnovo, Bulgaria e-mail: milenjh@yahoo.com

Isaac Iroegbu Ibenme Cairo wireless computer engineerng institute, Cairo, Egipt e-mail: prof ut jacob@earthling.net

Ikechukwu Clement Ijeh Cairo wireless computer engineering institute Cairo, Egipt e-mail: prof ut jacob@earthling.net

Bozhidar Z. Iliev Institute for Nuclear Research and Nuclear Energy 72 blvd.Tzarigradsko chaussee, 1784 Sofia, Bulgaria e-mail: bozho@inrne.bas.bg

Oladippo Daniel Ipadeola Cairo wireless computer engineering institute Cairo, Egipt e-mail: prof ut jacob@earthling.net

Angel Ivanov Technical University, 8 Kl. Ohridski blvd. 1000 Sofia, Bulgaria

Rossen Ivanov Institute for Nuclear Research and Nuclear Energy, 1784 Sofia, Bulgaria and School of Mathematics, Trinity College Dublin, Dublin 2, Ireland e-mail: ivanovr@tcd.ie

Branimir Kiradjiev University of forestry 10 Kliment Okhridsky Blvd. 1756 Sofia, Bulgaria e-mail: Kiradjiev2003@yahoo.com

Mihail Konstantinov University of Architecture, Civil Engeneering and Geodesy 1, Christo Smirnenski blvd., 1046 Sofia, Bulgaria e-mails: mmk_{−}fte@uacg.bg

Assen V. Kyuldjiev Institute for Nuclear Research and Nuclear Energy, 72 Tsarigradsko chaussee, 1784 Sofia, Bulgaria e-mails: kyuljiev@inrne.bas.bg

Nikolay A. Kostov Institute of Electronics, 72 Tsarigradsko chaussee, 1784 Sofia, Bulgaria e-mails: nakostov@ie.bas.bg

Julian Ławrynowicz Institute of Physics, University of Łódź Pomorska 149/153, PL-90-236 Łódź, Poland Institute of Mathematics, Polish Academy of Sciences Łódź Branch, Banacha 22, PL-90-238 Łódź, Poland e-mail: jlawryno@uni.lodz.pl

Sadahiro Maeda Department of Mathematics, Shimane University, Matsue 690-8504, Japan e-mail: smaeda@riko.shimane-u.ac.jp

Mancho Manev Department of Geometry, Faculty of Mathematics and Informatics, University of Plovdiv, 236 Bulgaria blvd., 4003 Plovdiv, Bulgaria e-mail: mmanev@yahoo.com

Vesna Manova-Erakovik Institute of Mathematics, Faculty of Natural Sciences and Mathematics St. Cyril and Methodius, University of Skopje, Macedonia e-mail: vesname@iunona.pmf.ukim.edu.mk

Marin Marinov New Bulgarian University, 21 Montevideo Str., 1618 Sofia, Bulgaria e-mail: mlmarinov@nbu.bg

Vanya Markova Institute of Control and System Research PO Box 565, 139 Russkii blvd, 4000 Plovdiv, Bulgaria e-mail: markovavanya@yahoo.com

Yasuo Matsushita Section of Mathematics, School of Engineering, University of Shiga Prefecture, Hikone 522-8533, Japan e-mail: yasuo@mech.usp.ac.jp

Hiroshi Matsuzoe Nagoya Institute of Technology, Gokiso-cho, Showa-ku, Nagoya, 466-8555, Japan e-mail: matsuzoe@nitech.ac.jp

Dimitar Mekerov Department of Geometry, Faculty of Mathematics and Informatics, University of Plovdiv, 236 Bulgaria blvd., 4003 Plovdiv, Bulgaria e-mail: mircho@pu.acad.bg

Velichka Milousheva Institute of Mathematics and Informatics, Acad. G. Bonchev Str. bl. 8, 1113 Sofia, Bulgaria e-mail: milousheva@vsu.bg

Galia Nakova University of Veliko Tarnovo, 2 T. Tarnovski str., 5000 Veliko Tarnovo, Bulgaria e-mail: gnakova@yahoo.com

Nikolay Nikolov Institute of Mathematics and Informatics Acad. G. Bonchev 8, 1113 Sofia, Bulgaria e-mails: nik@math.bas.bg

Peace Udom Nwaobilor Cairo wireless computer engineering institute Cairo, Egipt e-mail: prof ut jacob@earthling.net

Petko Petkov Department of Systems and Control, Technical University of Sofia, 1756 Sofia, Bulgaria, e-mails: php@tu-sofia.bg

Petar R. Popivanov Institute of mathematics and informatics, acad. G. Bontchev str., 1113 Sofia, Bulgaria e-mail: angp@math.bas.bg

Kouei Sekigawa Department of Mathematics, Faculty of Science, Niigata University, Niigata, 950-2181, Japan e-mail: sekigawa@sc.niigata-u.ac.jp

Slavka Slavova Dobrich Pedagogical College block 12 Dobrotica 9300 Dobrich, Bulgaria e-mail: slavova@dobrich.net

Petar Stoev University of Architecture, Civil Engeneering and Geodesy 1, Christo Smirnenski blvd. 1046 Sofia, Bulgaria

Tadashi Sugiyama Division of Mathematics and Mathematical Science, Department of Computer Science and Engineering, Nagoya Institute of Tecnhnology, Gokisocho Showaku Nagoya 466-8555 Japan e-mail: tadashi@zelus.ics.nitech.ac.jp

Marta Teofilova Department of Geometry, Faculty of Mathematics and Informatics, University of Plovdiv, Plovdiv 4003, Bulgaria e-mail: mar@gbg.bg

S. Udagawa Department of Mathematics, Shimane University, Matsue 690-8504, Japan

Tihomir Valchev Institute for Nuclear Research and Nuclear Energy 72 Tzarigradsko chaussee, 1784 Sofia, Bulgaria e-mail: valtchev@inrne.bas.bg

George Venkov Technical University, 8 Kl. Ohridski blvd. 1000 Sofia, Bulgaria

Ilona Zasada Solid State Physics Department, University of Łódź Pomorska 149/153, PL-90-236 Łódź, Poland e-mails: izasada@mvii.uni.lodz.pl